An investor invests $10000 in stocks trading at $100. If the price for these shares goes up by 30% and ignoring dividends, what would be the investor’s expected rate of return? Assuming the investor borrows another $10000 from a broker and invests in the same stocks to buy more of the same shares and the margin loan interest is 9% per year, what will his rate of return be now if the stocks goes up by 30%, ignoring dividends, again ? Suppose the investor only borrows $5000 at the same interest rate of 9% per year, what will the rate of return be if the share price goes up by 30%? If it goes down by 30%, and if it remains unchanged? 


1
Expert's answer
2021-02-16T11:23:26-0500

"Rate of return=ending value of investiment- multiply by 100%"

"=(10000*130)-(10000*100)\/(10000*10)*100%"

"=30%"

When share price goes up by 30%"=30*20000-900\/(20000*100)*100"

"=30%"

When share price goes up by 30%

"=30*1500-900\/(100*15000)*100"

"=30"

When share price reduces by 30%

"=-30*15000-900\/(100*15000)*100"

"=-30%"

When share price remains unchanged

"=0*15000-900\/(100*15000)*100"

"=-0.06"